You purchased 2,500 shares of the New Fund at a price of $20 per share at the beginning of the year. You paid a front-end load of 4%. The securities in which the fund invests increase in value by 11% during the year. The fund's expense ratio is 2.7%. What is your rate of return on the fund if you sell your shares at the end of the year? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Rate of return %
Answer #1

Opening value of investment= (2500*20)/(1-4%)

Opening value of investment=52,083.3333

________________________________________________

Closing value of investment= 50,000*(1+11%-2.7%)

Closing value of investment = 54150

______________________________

Rate of return= (closing value- opening value)/opening value

Rate of return = (54150-52088.3333)/52088.3333

Rate of return = 3.97%